OpenStudy (anonymous):

Robin is making bead necklaces from 60 beads. She wants each necklace to have the same number of beads. She determines she could make 10 necklaces with 6 beads or 2 necklaces with 30 beads. Which is not a possibility for her necklaces? * 4 necklaces of 15 beads 13 necklaces of 5 beads 3 necklaces of 20 beads 1 necklace with 60 beads

OpenStudy (anonymous):

help pls

OpenStudy (mathmath333):

4*15=60 13*5=65 3*20=60 60*1=60

OpenStudy (anonymous):

tyyaa

OpenStudy (anonymous):

ty,yasa